OpenStudy (anonymous):

A motorboat travels 118km in 2 hours going upstream. It travels190 going downstream in the same amount of time. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (lgbasallote):

upstream speed = speed of motorboat (in still water) - current downstream speed = speed of motorboat (in still water) + current let b = speed of boat in still; c = current you are given a distance and some time...you're looking for the rate of boat in still and rate of current... Hint: first thing you'll do here is to solve for the downstream speed and the upstream speed..do you know how?

OpenStudy (lgbasallote):

to find downstream speed...you use rate = \(\large \frac{distance}{time}\)

OpenStudy (dumbcow):

side note: i have seen so many of these problems, they need to come up with different rate problems

OpenStudy (anonymous):

thnx Igbasallote, that helped

OpenStudy (lgbasallote):

well..i'll continue anyway :DDD so downstream speed = \(\Large \frac{190}{2}\) = 95 km/h upstream speed = \(\Large \frac{118}{2}\) = 59 km/h we put this back in the system... 95 km/h = b + c <----where b is boat speed; c is current speed 59 km/h = b - c we can use substitution method here by letting c = 95 - b..we put htat in the second equation... 59 = b -(95 - b) 59 = b -95 + b 95 + 59 = 2b 154 = 2b b = 77 km/h knowing this... we can find c.. 95 = 77 + c 95 - 77 = c 18 km/h = c

OpenStudy (anonymous):

wow ur good. thnx!

OpenStudy (lgbasallote):

heh thanks ^_^ <tips hat> hope that explanation was clear enough :DDD

OpenStudy (anonymous):

does (-2v^2 +7v-6)+(v^2 +5v+4) equal -v^2 +12v-2?

OpenStudy (lgbasallote):

\(\Huge \mathbb{YEP!}\) \(\Huge \checkmark\)

OpenStudy (anonymous):

lol. ty

OpenStudy (anonymous):

\[\sqrt{-40}/\sqrt{-2}=? and \sqrt{-5}*\sqrt{-20}=?\]

OpenStudy (lgbasallote):

\(\huge \frac{\sqrt{40}}{\sqrt{-2}}\) = \(\huge \sqrt{\frac{40}{-2}}\) as for multiplication... \(\huge \sqrt{-5} * \sqrt{-20} = \sqrt{(-5)(-20)}\) does that give you any hints?

OpenStudy (anonymous):

rectangle length is 5 times twice its width, and the area is88ft^2. wat are the length and width of the rectangle?

OpenStudy (lgbasallote):

let L = length; w = width L = 5(2w) a = 88 a = lw a = 5(2w)(w) substitute values..

OpenStudy (anonymous):

gtg